# GenTKG: Generative Forecasting on Temporal Knowledge Graph

### v3

@@ Abstract
The rapid advancements in large language models (LLMs) have ignited interest in the temporal knowledge graph (tKG) domain, where conventional embedding-based and rule-based methods dominate. The question remains open of whether pre-trained LLMs can understand structured temporal relational data and replace them as the foundation model for temporal relational forecasting. Therefore, we bring temporal knowledge forecasting into the generative setting. However, challenges occur in the huge chasms between complex temporal graph data structure and sequential natural expressions LLMs can handle, and between the enormous data sizes of tKGs and heavy computation costs of finetuning LLMs. To address these challenges, we propose a novel retrieval-augmented generation framework named GenTKG combining a temporal logical rule-based retrieval strategy and few-shot parameter-efficient instruction tuning to solve the above challenges, respectively. Extensive experiments have shown that GenTKG outperforms conventional methods of temporal relational forecasting with low computation resources using extremely limited training data as few as 16 samples. GenTKG also highlights remarkable cross-domain generalizability with outperforming performance on unseen datasets without re-training, and in-domain generalizability regardless of time split in the same dataset. Our work reveals the huge potential of LLMs in the tKG domain and opens a new frontier for generative forecasting on tKGs.  [^1] Under review. The code and data will be released soon.
@@
[…]
@@ 1 Introduction
Conventional embedding-based graph representation learning methods (Goel et al., 2020; Han et al., 2020a; Sun et al., 2021; Yang et al., 2020) require carefully designed models that embed indexed quadruples into hidden latent space and hence lose the semantic aspects of events in tKGs. Besides, they require separate training for different datasets and hence suffer to handle even slight dataset modification and time split adaptation. In stark contrast, the rule-based methods Liu et al. (2022) focus on mining temporal logic rules within the tKG graph structure in a symbolic way using neural networks. However, it posesses limited scalability to only similar datasets sharing silimar rules. With the huge advancements emerging with numerous large language models (LLMs) (Wei et al., 2022), for example utilizing the emergent in-context learning (ICL) ability of LLMs (Dong et al., 2022) by sequentializing temporal ascending ordered tKG facts to texts but failed to compete with the above conventional methods (Lee et al., 2023). The question remains open: Can pre-trained LLMs understand structured temporal relational data and replace conventional methods as the foundation model for temporal relational forecasting?
@@
Our approach offers a foundational framework for future explorations in generative forecasting on temporal knowledge graphs. Our contributions can be summarized as follows:
@@ 4.1 Main Results
Compared to embedding-based method. For all datasets, GenTKG outperforms its best embedding-based model xERTE on ICEWS14, ICEWS18, GDELT, and Timetraveler on YAGO. Specifically, the highest performance gain is observed on GDELT with more than 58% higher on Hits@1.
@@
Compared to in-context-learning method. We analyze the performance of GenTKG on different Language Model instantiations, i.e. GPT-NeoX-20B and LLaMA2-7B respectively. For GPT-NeoX-20B, we apply only the first retrieval phase of GenTKG due to hardware limitations. However, huge performance increase is observed for all three metrics on all datasets even with pure retrieval-augmented in-context learning. For LLaMA2-7B, the performance gain of Hits@1 has increased remarkably even outperforming GPT-NeoX-20B which has two times more parameters, indicating the potential for greater performance of our proposed GenTKG framework if applied to larger language models.
@@ 4.2 Cross-domain Generalization
To answer the second question of GenTKG’s performance in the inductive setting, the empirical results indicate that the GenTKG framework manifests a substantial capability for cross-dataset generalization. Specifically, once the LLM has been aligned to the tKG forecasting task in the second phase on any dataset, the LLM can be applied directly to any other datasets. Therefore, on a new dataset, GenTKG only requires dataset-specific temporal-logical rule-based retrieval to formulate proper prompts from the first phase, and can directly infer the predictions without retraining in the second phase. As shown in Figure 3(a), all methods are trained and evaluated on GDELT, except that the LLM in inductive GenTKG is trained ICEWS14. Still, the inductive GenTKG delivers comparable performance metrics on GDELT to conventional methods with a minor performance drop compared to the orginal trained GenTKG. We further demonstrate similar inductive results by cross-checking the training and evaluation datasets as shown in Figure 3(b). Although the LLM is trained exclusively on one dataset, it still delivers comparable metrics on disparate datasets, closely approximating the outcomes of methods that were trained specifically on the identical evaluation dataset. This notable characteristic implies that the GenTKG framework is effectively capturing the underlying task-related features, as opposed to merely carefully-designed for the dataset data, a limitation commonly shared in conventional methods.
Figure 3: Cross-Domain Inductive Setting. (a) Single dataset evaluation. All training and evaluation is on GDELT except inductive GenTKG is trained on ICEWS14. (b) Cross-checking. We cross-check the trained LLaMA2 in GenTKG on different training datasets and evaluation datasets. The performance drop compared to the original training setting takes up only small percentages. Even higher performance than ICL can be observed. More discussions about experiment settings and analysis are given in Appendix B.2, explaining the huge relative difference on GDELT is due to its poor baseline performances.
@@ 4.4 Ablation study
We undertake the ablation studies on ICEWS14 to evaluate the contribution of each phase in GenTKG with three distinct variants of the GenTKG: TLR, FIT, and TLR+FIT configurations. Here, TLR represents the variant that exclusively employs temporal logical rule-based retrieval on top of ICL learning, FIT denotes the variant solely implementing few-shot parameter-efficient instruction tuning with naive fact retrieval (Lee et al., 2023), and TLR+FIT encapsulates the integration of all components within GenTKG. Figure 5(a) draws the conclusion that both phases in GenTKG framework contribute to distinct performance improvements. The whole pipeline enables GenTKG the ability to outperform existing methods.
@@ 5 Discussion
Q4: How does history length affect ICL’s performance?
@@ 7 Limitation and Future Directions
GenTKG is tailored to generative forecasting on temporal knowledge graph and can be applied to a wide variety of downstream tasks with generative forecasting setting, such as recommendation system, anomly detection, etc. It can also power search and serve to improve users’ lives. GenTKG can help protect data with its generalizability which requires less training over various datasets. The risk of GenTKG might comes from risks inherited in open-source LLMs, such as hallucinations.
@@
The datasets used in this research work is open-sourced and can be seen on references. We derive some datasets from the original version within the intended use term. The code and source will be released after review.
@@ Appendix A Related Works
Temporal knowledge graphs (tKGs) are multi-relational, directed graphs with labeled timestamped edges between entities (nodes). Let $\mathcal{E}$ and $\mathcal{P}$ represent a finite set of entities and predicates. A quadruple $(e_{s},r,e_{o},t)$ represents a timestamped and labeled edge between a subject entity $e_{s}\in\mathcal{E}$ and an object entity $e_{o}\in\mathcal{E}$ at a timestamp $t\in\mathcal{T}$ Let $\mathcal{F}$ represent the set of all true quadruples, i.e., real events in the world, the temporal knowledge graph forecasting is the task of predicting missing object entity at timestamp $t$, i.e. $(e_{s},r,?,t)$ based on a set of observed facts $\mathcal{O}$ before $t$, which is a subset of $\mathcal{F}$. Current methods can be categorized into two streams. On the one hand, embedding-based models learn representations of the quadruples with carefully designed embedding modelsHan et al. (2020a); Goel et al. (2020); Sun et al. (2021); Han et al. (2020b); Ding et al. (2022). On the other hand, the rule-based methods mine the temporal logical rules extracted and extract candidates directly on the temporal knowledge graphsLiu et al. (2022).
@@
The semantic part stored in the temporal knowledge graphs is heavily overlooked in either embedding-based or rule-based temporal knowledge graph methods. Early explorers had tryouts in introducing language models in the TKG domain, some fused pre-trained language representations to the temporal knowledge embeddings (Han et al., 2022), and some flattened explicit temporal events with the emergent in-context learning ability of large language models however not comparable with conventional performance (Lee et al., 2023).
@@ Appendix B Supplimentary Materials
B.1 Implementation details.
Experiment hyperparameters will be released in code soon. We run experiments 3 times and take averages with A40 GPU.
B.2 Discussion on Cross-domain Generalizability
@@
To assess the cross-domain generalizability according to the 4 test benchmarks in this paper, we conduct 4 series of cross-domain inductive settings respective to each benchmark. We define that a series of cross-checking settings consists of a center evaluation dataset $\mathcal{A}$ with the other three cross-checking datasets denoted as $\mathcal{B,C,D}$. Inside a series, a single evaluation on $\mathcal{A}$ is conducted by comparing all inference results on the center $\mathcal{A}$ including (1) all baseline methods trained on $\mathcal{A}$, (2) original GenTKG trained on $\mathcal{A}$, and (3) generalized GenTKG trained one of the other three cross-checking dataset, e.g. $\mathcal{B}$. In total, $4\times 3=16$ cross-checking experiments are conducted.
@@ 11             $\mathcal{G}_{s}(e_{s},r,t)\leftarrow\mathcal{G}_{s}(e_{s},r,t)\cup fact$
B.4 Dataset Statistics

### v5

@@ Abstract
The rapid advancements in large language models (LLMs) have ignited interest in the temporal knowledge graph (tKG) domain, where conventional embedding-based and rule-based methods dominate. The question remains open of whether pre-trained LLMs can understand structured temporal relational data and replace them as the foundation model for temporal relational forecasting. Therefore, we bring temporal knowledge forecasting into the generative setting. However, challenges occur in the huge chasms between complex temporal graph data structure and sequential natural expressions LLMs can handle, and between the enormous data sizes of tKGs and heavy computation costs of finetuning LLMs. To address these challenges, we propose a novel retrieval-augmented generation framework named GenTKG combining a temporal logical rule-based retrieval strategy and few-shot parameter-efficient instruction tuning to solve the above challenges, respectively. Extensive experiments have shown that GenTKG outperforms conventional methods of temporal relational forecasting with low computation resources using extremely limited training data as few as 16 samples. GenTKG also highlights remarkable cross-domain generalizability with outperforming performance on unseen datasets without re-training, and in-domain generalizability regardless of time split in the same dataset. Our work reveals the huge potential of LLMs in the tKG domain and opens a new frontier for generative forecasting on tKGs. The code and data are released here: https://github.com/mayhugotong/GenTKG.
@@
with Large Language Models
Ruotong Liao1,2,   Xu Jia3,   Yangzhe Li3,   Yunpu Ma1,2,4,   Volker Tresp1,2 1LMU Munich     2Munich Center for Machine Learning (MCML) 3Technical University of Munich   4Siemens AG ruotong.liao@outlook.com,   cognitive.yunpu@gmail.com volker.tresp@lmu.de
@@ 1 Introduction
In tKG, the first embedding-based representation learning method is introduced by (Ma et al., 2019). The following conventional embedding-based methods (Goel et al., 2020; Han et al., 2020a; Sun et al., 2021; Yang et al., 2020; Li et al., 2022) require carefully designed models that embed indexed quadruples into hidden latent space and hence lose the semantic aspects of events in tKGs. Besides, they require separate training for different datasets and hence suffer to handle even slight dataset modification and time split adaptation. In stark contrast, the rule-based methods Liu et al. (2022) focus on mining temporal logic rules within the tKG graph structure in a symbolic way using neural networks. However, it possesses limited scalability to only similar datasets sharing similar rules. With the huge advancements emerging with numerous large language models (LLMs) (Wei et al., 2022), for example utilizing the emergent in-context learning (ICL) ability of LLMs (Dong et al., 2022) by sequentializing temporal ascending ordered tKG facts to texts but failed to compete with the above conventional methods (Lee et al., 2023). The question remains open: Can pre-trained LLMs understand structured temporal relational data and replace conventional methods as the foundation model for temporal relational forecasting?
@@
Our approach offers a foundational framework for future explorations in generative forecasting on temporal knowledge graphs. Our contributions are:
@@ 4.1 Main Results
Compared to embedding-based methods. For all datasets, GenTKG outperforms its best embedding-based model xERTE on ICEWS14, ICEWS18, GDELT, and Timetraveler on YAGO. Specifically, the highest performance gain is observed on GDELT with more than 58% higher on Hits@1.
@@
Compared to in-context-learning method. We analyze the performance of GenTKG on different Language Model instantiations, i.e. GPT-NeoX-20B and LLaMA2-7B respectively. For GPT-NeoX-20B, we apply only the first retrieval phase of GenTKG due to hardware limitations. However, a huge performance increase is observed for all three metrics on all datasets even with pure retrieval-augmented in-context learning. For LLaMA2-7B, the performance gain of Hits@1 has increased remarkably even outperforming GPT-NeoX-20B which has two times more parameters, indicating the potential for greater performance of our proposed GenTKG framework if applied to larger language models.
@@ 4.2 Cross-domain Generalization
To answer the second question of GenTKG’s performance in the generalization setting, the empirical results indicate that the GenTKG framework manifests a substantial capability for cross-dataset generalization. Specifically, once the LLM has been aligned to the tKG forecasting task in the second phase on any dataset, the LLM can be applied directly to any other dataset. Therefore, on a new dataset, GenTKG only requires dataset-specific temporal-logical rule-based retrieval to formulate proper prompts from the first phase, and can directly infer the predictions without retraining in the second phase. As shown in Figure 3(a), all methods are trained and evaluated on GDELT, except that the LLM in generalized GenTKG is trained ICEWS14. Still, the generalized GenTKG delivers comparable performance metrics on GDELT to conventional methods with a minor performance drop compared to the original trained GenTKG. We further demonstrate similar generalization results by cross-checking the training and evaluation datasets as shown in Figure 3(b). Although the LLM is trained exclusively on one dataset, it still delivers comparable metrics on disparate datasets, closely approximating the outcomes of methods that were trained specifically on the identical evaluation dataset. This notable characteristic implies that the GenTKG is effectively capturing the underlying task-related features, as opposed to merely carefully designed for the dataset data, a limitation commonly shared in conventional methods.
Figure 3: Cross-Domain Generalization Setting. (a) Single dataset evaluation. All training and evaluation is on GDELT except generalized GenTKG, which is trained on ICEWS14. (b) Cross-checking. We cross-check the trained LLaMA2 in GenTKG on different training datasets and evaluation datasets. The performance drop compared to the original training setting takes up only small percentages. Even higher performance than ICL can be observed. More discussions about experiment settings and analysis are given in Appendix B.1, explaining the huge relative difference on GDELT is due to its poor baseline performances.
@@ 4.4 Ablation study
We undertake ablation studies on ICEWS14 to evaluate the contribution of each phase in GenTKG with three distinct variants of the GenTKG: TLR, FIT, and TLR+FIT configurations. Here, TLR represents the variant that exclusively employs temporal logical rule-based retrieval on top of ICL learning, FIT denotes the variant solely implementing few-shot parameter-efficient instruction tuning with naive fact retrieval (Lee et al., 2023), and TLR+FIT encapsulates the integration of all components within GenTKG. Figure 5(a) draws the conclusion that both phases in GenTKG framework contribute to distinct performance improvements. The whole pipeline enables GenTKG the ability to outperform existing methods.
@@ 5 Discussion
Q4: How does history length affect GenTKG’s performance?
@@ 7 Limitation and Future Directions
GenTKG is tailored to generative forecasting on temporal knowledge graphs and can be applied to a wide variety of downstream tasks with generative forecasting settings, such as recommendation systems, anomaly detection, etc. It can also power search and serve to improve users’ lives. GenTKG can help protect data with its generalizability which requires less training over various datasets. The risk of GenTKG might come from risks inherited in open-source LLMs, such as hallucinations.
@@
The datasets used in this research work are open-sourced and can be seen in references. We derive some datasets from the original version within the intended use term. For the GDELT dataset, as stated in the terms of use of GDELT [^2] https://www.gdeltproject.org/about.html#termsofuse, this project is an open platform for research and analysis of global society and thus all datasets released by the GDELT Project are available for unlimited and unrestricted use for any academic, commercial, or governmental use of any kind without fee. One may redistribute, rehost, republish, and mirror any of the GDELT datasets in any form. However, any use or redistribution of the data must include a citation to the GDELT Project and a link to this website (https://www.gdeltproject.org/). ICEWS follows the MIT license on its official website (https://github.com/andybega/icews?tab=MIT-2-ov-file) and YAGO is licensed under a Creative Commons Attribution 4.0 International License (https://yago-knowledge.org/).
Acknowledgements
This work was funded by the Munich Center for Machine Learning and supported by the Federal Ministry of Education and Research and the State of Bavaria. We are thankful for the helpful discussions with Yushan Liu from Siemens AG and Zifeng Ding from LMU Munich.
@@ Appendix A Related Works
Temporal knowledge graphs (tKGs) are multi-relational, directed graphs with labeled timestamped edges between entities (nodes). Let $\mathcal{E}$ and $\mathcal{P}$ represent a finite set of entities and predicates. A quadruple $(e_{s},r,e_{o},t)$ represents a timestamped and labeled edge between a subject entity $e_{s}\in\mathcal{E}$ and an object entity $e_{o}\in\mathcal{E}$ at a timestamp $t\in\mathcal{T}$ Let $\mathcal{F}$ represent the set of all true quadruples, i.e., real events in the world, the temporal knowledge graph forecasting task predicts missing object entity at timestamp $t$, i.e. $(e_{s},r,?,t)$ based on a set of observed facts $\mathcal{O}$ before $t$, which is a subset of $\mathcal{F}$. Current methods can be categorized into two streams. Embedding-based models learn representations of the quadruples with carefully designed embedding models Han et al. (2020a); Goel et al. (2020); Sun et al. (2021); Han et al. (2020b); Ding et al. (2022). Rule-based methods mine the temporal logical rules extracted and extract candidates directly on the tKGs Liu et al. (2022).
@@
The semantic part stored in the temporal knowledge graphs is heavily overlooked in either embedding-based or rule-based temporal knowledge graph methods. Early explorers had tryouts in introducing language models in the TKG domain, some fused pre-trained language representations to the temporal knowledge embeddings (Han et al., 2022), and some flattened explicit temporal events with the emergent in-context learning ability of large language models however not comparable with conventional performance (Lee et al., 2023).  (Ding et al., 2023) explores LLM in the zero-shot relational learning settings in the TKG forecasting task.
@@ Appendix B Supplimentary Materials
B.1 Discussion on Cross-domain Generalizability
@@
To assess the cross-domain generalizability according to the 4 test benchmarks in this paper, we conduct 4 series of cross-domain generalization settings respective to each benchmark. We define that a series of cross-checking settings consists of a center evaluation dataset $\mathcal{A}$ with the other three cross-checking datasets denoted as $\mathcal{B,C,D}$. Inside a series, a single evaluation on $\mathcal{A}$ is conducted by comparing all inference results on the center $\mathcal{A}$ including (1) all baseline methods trained on $\mathcal{A}$, (2) original GenTKG trained on $\mathcal{A}$, and (3) generalized GenTKG trained one of the other three cross-checking dataset, e.g. $\mathcal{B}$. In total, $4\times 3=16$ cross-checking experiments are conducted.
@@ 12             $\mathcal{G}_{s}(e_{s},r,t)\leftarrow\mathcal{G}_{s}(e_{s},r,t)\cup fact$
B.3 Implementation Details
We run experiments 3 times and take averages with A40 GPU. For the TLR part, we use the rule length of 1, the number of random walks of 200, the time window of the maximum length of each dataset, and the maximum history length of 50. In the FIT training, we use the batch size of 1024, the learning rate of $43e-$ 4, the context length of 4096, the target length of 128, the LoRA rank of 8, the LoRA dropout rate of 0.05, and few-shot tuning of 1024-shots. Besides, we use the Adam optimizer (Kingma and Ba, 2014).
B.4 Supplementary Statistics